OpenStudy (anonymous):

Evaluate the function rule for the given value. f(x) 5^x for x = 2 a. 25 b. 10 c. 32 d. 5

OpenStudy (anonymous):

given f(x)=5^x so,f(2)=5^2 =25

OpenStudy (anonymous):

thank you , i get this one... its actually quit easy lol but thanks

OpenStudy (anonymous):

welcome =)